Status Report

NOAA GOES SATOPS Morning Report 19 Jan 2001

By SpaceRef Editor
January 19, 2001
Filed under ,

SATOPS MORNING REPORT – January 19, 2001

NWS declares Critical Weather Days for 019/1200Z – 022/1200Z.

POES

No change in the status of the POES spacecraft.

From 1741Z to 1804Z on January 18:  Fairbanks power failure. 
Switch gear controlling all breakers failed.

NOAA-14 rev 31203 / F at 1801Z on January 18:  Not acquired due
to power failure at Fairbanks.  Load was delayed to rev 31204W. 
Data lost:  L1B, L3A, L5A.

NOAA-14 rev 31204 / W at 1927Z on January 18:  Load successfully
uplinked.  G1A fully recovered from rev 31203.

NOAA-14 rev 31207 / F at 0037Z on January 19:  In addition to the
scheduled data, we recovered 100% of G5B, which was originally scheduled
for recovery during rev 31203.

NOAA-16 revs 1682 & 1684 on January 18:  CPU-A anomaly resolution-related
commanding continues.

NOAA-16 rev 1688 at 0304Z on January 19:  SBUV activation sequence
begins from the stored command table.

DMSP

No change in the status of the DMSP spacecraft.

DMSP operations were nominal over the past 24 hours.

GOES

No change in the status of the GOES spacecraft.

At 0446Z on January 19:  GOES-8 image frame break due to loss of
signal strength.

At 1833Z, 1915Z, and 1924Z on January 18:  GOES-10 image frame
breaks due to high bit error rate.

 

Questions or comments to [email protected]

SpaceRef staff editor.